OpenStudy (anonymous):

A wind turbine has blades that sweep an area of 2000 m2. It converts the power available in the wind to electrical power with an efficiency of 50%. What is the electrical power generated if the wind speed is 10 m s–1? (The density of air is 1.3 kg m–3.)

OpenStudy (anonymous):

Find kinetic energy of the blade per unit time = power input = 1/2 m v^2 /t They have given volume density and area, which would give mass/ t 2000*10*1.3 = 26000J --------Power input = 1/2 * 26000 * (10)^2 = 1300kW efficiency = power output/power input power output = 0.5*1300kw = 650kw